OpenStudy (anonymous):

After a concert for 1,600 people, only 8 people said they would not go to see the band again. What percent of the people who went to the concert said they would not go to see the band again? 200 50% 8% 0.5% Please help me I will give you a medal and I will become a fan. Thank you!

OpenStudy (anonymous):

@texaschic101

OpenStudy (anonymous):

@texaschic101

OpenStudy (anonymous):

last one

OpenStudy (texaschic101):

basically, your problem is : 8 is what percent of 1600 percentage formula : you can use this with any percent problem is/of = %/100 is = 8 % = x of = 1600 now we sub 8/1600 = x/100 cross multiply 1600x = 8(100) 1600x = 800 x = 800/1600 x = 0.5 %

OpenStudy (anonymous):

Thank you so much!!!!

OpenStudy (texaschic101):

anytime...happy to help :)

OpenStudy (anonymous):

Dang your smart xD

OpenStudy (texaschic101):

not on some things.....geometry is a problem for me

OpenStudy (anonymous):

Omg thanks I got 100% :D

OpenStudy (texaschic101):

sweet....time for the happy dance :)